OpenStudy (anonymous):

WILL GIVE MEDAL: Sally invests $4000 in a savings account which compounds interest monthly. After 10 years, she has a balance of $5,451.52. What is the annual interest rate of the account, rounded to the nearest 0.01%?

OpenStudy (anonymous):

I have it set up just need help solving it. :o

OpenStudy (anonymous):

n = 10*c= 12*10 = 120 I = rate/c = ? P = 4000 A = 5451.20

OpenStudy (anonymous):

I meant 5451.52

OpenStudy (anonymous):

This is what I have it set up as: 5451.52= 4000(1+ r/12)^ 12*10

OpenStudy (anonymous):

good.

OpenStudy (anonymous):

I just dunno how to solve for the r I've been lost :o

OpenStudy (anonymous):

\[1.36288 = (1+\frac{r}{12})^{120}\\ \sqrt[120]{1.36288} = (1+ \frac{r}{12})\]

OpenStudy (anonymous):

\[1.00258 = ( 1 + \frac{r}{12})\\0.0025833 = \frac{r}{12}\]

OpenStudy (anonymous):

Awesome! So after the last step do I multiply by 12 to get the r?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got .0309996 did yo happen to get the same answer?

OpenStudy (anonymous):

X 100 = ?

OpenStudy (anonymous):

3.10% is..that an intuitive answer? would that be a reasonable interest rate for this financial interaction?

OpenStudy (anonymous):

Yes, it would! Thank you so much for your help! I appreciate it! :)